# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

### v1

@@ Abstract
Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime. This paper introduces a novel, integrated framework that leverages the power of transformer neural networks and deep reinforcement learning (DRL) algorithms to optimize maintenance actions. Our approach employs the transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the Remaining Useful Life (RUL) of equipment. Simultaneously, the DRL component of our framework provides cost-effective and timely maintenance recommendations. We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions. Consequently, our pioneering approach provides an innovative data-driven methodology for prescriptive maintenance, addressing key challenges in industrial operations and leading the way to more efficient, cost-effective, and reliable systems.
@@ I Introduction
The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies. With these advancements, however, comes the heightened need for reliable, efficient, and cost-effective maintenance systems. In particular, as manufacturing processes become increasingly automated, the costs associated with machine downtime rise substantially. Traditionally, human operators have been required to monitor machinery constantly. The COVID-19 pandemic has complicated this scenario, elevating the costs associated with labor-intensive activities and emphasizing the need for reducing on-site staff to minimize infection risks.
Prescriptive maintenance has emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [1]. It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes. According to ISO standard 13381-1 [2], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning.
Current predictive maintenance systems largely rely on heuristic models or conventional machine learning algorithms. While effective to some extent, these approaches often fail to capture complex temporal dependencies in sensor data, limiting their ability to accurately predict RUL or prescribe optimal maintenance actions. Furthermore, many existing solutions require extensive labeled datasets, which are costly and time-consuming to generate.
In response to these challenges, this paper proposes a novel, data-driven framework that leverages the advanced capabilities of Transformer neural networks and DRL algorithms. Transformers excel at modeling long-term dependencies in sequential data and are highly adaptable, making them ideal for handling complex industrial sensor data. DRL algorithms, on the other hand, excel at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks.
The framework aims to achieve the following:
  • Accurately predict the RUL of machinery using Transformer-based models, thus providing early warnings for potential failures.
  • Employ DRL algorithms to offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs.
Unlike existing solutions that address either predictive or prescriptive maintenance, our framework integrates both aspects seamlessly. It leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations. By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges. Additionally, we validate our approach using the NASA C-MPASS dataset, achieving superior performance in both predicting RUL and optimizing maintenance actions, even when compared to traditional neural networks on larger datasets.
Contributions:
  • We are the first to address the unique challenges of limited dataset sizes in the prescriptive maintenance landscape.
@@
  • We propose a DRL-based recommendation system for maintenance actions, reducing both labor involvement and operational costs.
@@
Organization. The rest of the paper is organized as follows: Section II reviews related work in the field of prescriptive maintenance and machine learning algorithms. Section III describes the proposed framework in detail. Section IV presents the experimental setup and results. Finally, Section V concludes the paper and outlines future directions for research.
@@ II Related Work
Corrective Maintenance is “fault-based maintenance” and is an unplanned mode of maintenance carried out only when equipment malfunctions are detected. Preventive Maintenance, or scheduled inspection and maintenance, generally involves halting the production line according to a predetermined schedule, with an emphasis on repairs before a failure occurs. These two types of maintenance methods are traditionally the most commonly adopted in the industry. However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance. Predictive Maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as Condition-Based Maintenance (CBM). Prescriptive Maintenance is generally considered an extension of Predictive Maintenance, optimizing maintenance strategies based on fault diagnosis results. It emphasizes identifying the inherent reasons for the occurrence of faults and subsequently directing the generation of operational and maintenance strategies [4]. Compared to traditional maintenance methods, Prescriptive Maintenance leverages historical system status to comprehensively forecast future system operations and optimize the resources required for maintenance, aiming to prevent waste from excessive maintenance or production losses due to malfunctions and downtime.
@@ II-A Deep Learning-Based Remaining Useful Life Prediction
Accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations. While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [5, 6]. Notably, they also entail labor-intensive manual feature engineering.
However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data. Deep learning, catalyzed by computational advances, emerges as a remedy. Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [7, 8]. Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
Amongst the deep learning architectures, Transformers, originally conceived for language tasks [9], have recently demonstrated efficacy in time-series data processing. For instance, Zhou et al. [10] integrated a time-embedding layer, augmenting the Transformer’s time-series prediction. Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
@@ II-B Deep Reinforcement Learning enabled Predictive Maintenance
DRL’s foray into predictive maintenance has primarily focused on refining maintenance schedules [11, 1, 12]. While these applications exhibit DRL’s prowess in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, eschewing actionable preventative measures. Gordon et al. [13] did venture into employing DRL for prescriptive maintenance, but the holistic integration of prediction and actionable recommendations remained elusive.
In the annals of maintenance research, preventive and predictive methods have been extensively probed, as highlighted by Ansari [14] and Nemeth [15]. Yet, the burgeoning domain of prescriptive maintenance, transcending mere forecasts to offer optimized interventions, yearns for a unified solution. While Ansari et al. [14] and Nemeth et al. [15] have illuminated this space, the prevalent trend has either skewed towards prediction or prescription. Despite separate advancements in Transformer-based predictions and DRL prescriptive models, a comprehensive framework synergizing these modalities remains absent. This research endeavors to traverse this void, proposing an integrated approach for prescriptive maintenance. By amalgamating the Transformer’s predictive capabilities with DRL’s prescriptive strengths, our framework envisions precise malfunction forecasts paired with optimized action plans.
@@ III Problem Formulation
In the modern industrial ecosystem, machinery and equipment form the backbone of operations. But with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability. The cascade effect is not just limited to production loss; it reverberates through aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [16]. Therefore, the primary challenges reside in not just predicting when machinery might fail but also understanding the myriad of preventive actions that can be employed effectively.
@@ III-A Predicting Remaining Useful Life (RUL)
At the core of predictive maintenance lies the notion of RUL, which provides an estimation of the duration a machine or its components will continue to function before encountering a failure. An accurate RUL prediction empowers industries, enabling them to strategically schedule maintenance activities and, in doing so, avert the unpredicted halting of operations. The RUL offers a dynamic and predictive window into the operational health of a machine. However, deducing this metric isn’t straightforward. The operational data of a machinery component is typically represented as a time series $X=\{x_{1},x_{2},\cdots,x_{t}\}$, where each $x_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions. The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
  • $T$ is the total expected operational lifespan of the machine or component, which might vary based on its usage, environmental factors, and maintenance history.
@@
The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors. The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $X$ to provide an accurate approximation of $R(t)$: $M(X)\approx R(t)$.
III-B Optimal Maintenance Recommendations with DRL
While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions. Upon gaining insights into potential machine downtimes, the challenge shifts to determining the most optimal maintenance action. Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
Consider a machinery component’s state $s$, encapsulating its health, RUL, and other relevant metrics. From this state, we want to choose an action $a$ from a set of potential actions $A$ to maximize the expected future reward. Given a machinery state $s$ which integrates various metrics (like its RUL, health indicators, operational constraints, etc.), the optimal maintenance action $a$ needs to be chosen from a pool of potential actions $A$. The essence of this problem is rooted in the principle of maximizing long-term rewards.
We can mathematically denote a value function $Q(s,a)$, which illustrates the expected cumulative benefit or reward of adopting action $a$ in state $s$. The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in A}Q(s,a)$, where
@@ III-B Optimal Maintenance Recommendations with DRL
  • The cumulative reward integrates both tangible (like cost savings) and intangible (like averting potential future issues) benefits.
The central challenge here is not only defining the rewards but also ensuring that the decision-making process is robust and can adapt to evolving machinery states and operational requirements.
@@ IV Proposed Framework
Our proposed framework consists of two primary components: a Transformer model for RUL prediction and a subsequent module for recommending maintenance actions. Our proposed framework can be illustrated in the following Algorithm 1 :
Input: Input sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$

### v2

@@ Abstract
Industrial systems demand reliable predictive maintenance strategies to enhance operational efficiency and reduce downtime. This paper introduces an integrated framework that leverages the capabilities of the Transformer model-based neural networks and deep reinforcement learning (DRL) algorithms to optimize system maintenance actions. Our approach employs the Transformer model to effectively capture complex temporal patterns in sensor data, thereby accurately predicting the remaining useful life (RUL) of an equipment. Additionally, the DRL component of our framework provides cost-effective and timely maintenance recommendations. We validate the efficacy of our framework on the NASA C-MPASS dataset, where it demonstrates significant advancements in both RUL prediction accuracy and the optimization of maintenance actions, compared to the other prevalent machine learning-based methods. Our proposed approach provides an innovative data-driven framework for industry machine systems, accurately forecasting equipment lifespans and optimizing maintenance schedules, thereby reducing downtime and cutting costs.
@@ I Introduction
The fourth industrial revolution, colloquially known as Industry 4.0, has ushered in unprecedented levels of automation and data exchange in manufacturing technologies [1]. With these advancements, however, comes the increased requests for reliable, efficient, and cost-effective maintenance systems. Prescriptive maintenance has thus emerged as a vital area of research and industrial application, offering the promise of proactive fault detection and optimized action planning [2]. It differs from reactive maintenance, which only kicks in post-failure, causing often unplanned and expensive downtimes. According to ISO standard 13381-1 [3], the predictive maintenance cycle includes data preprocessing, failure prognostics, and action planning. However, current predictive maintenance systems largely rely on heuristic models for action planning or conventional machine learning algorithms for sensor data processing. While effective to some extent, existing approaches often fail to capture complex temporal dependencies in sensor data due to issues like vanishing or exploding gradients, limiting their ability to accurately predict Remaining Useful Life (RUL).
Apart from that, there is a trend in recent works to leverage Reinforcement Learning (RL) to make plans in maintenance tasks [4]. RL excels at making optimal decisions in complex environments, offering a powerful tool for action planning in maintenance tasks. As a result, the RL-based decisions offer optimal maintenance action schedules that balance the trade-off between equipment reliability and maintenance costs. However, these methods consistently encounter the cold start problem [5], coupled with low learning efficiency. In some special cases such as aeroengine, these methods also suffer from a lack of reliability and explainability. This often leads to their decisions being perceived as untrustworthy by users.
Unlike existing solutions that address either predictive or prescriptive maintenance, we propose a framework named TranDRL, which integrates both aspects seamlessly. TranDRL leverages advanced machine learning techniques to not only accurately predict RUL but also to convert these predictions into actionable maintenance recommendations. By integrating these two critical components, our framework provides a comprehensive solution for industrial maintenance challenges by Transformer neural networks and Reinforcement Learning with Human Feedback (RLHF) algorithms. Transformers excel at modeling long-term dependencies in sequential data by the attention mechanism and are highly adaptable, making them ideal for handling complex industrial sensor data. By leveraging the expertise and insights of human operators or maintenance professionals, the algorithm can efficiently guide its actions, accelerating the convergence of the DRL algorithm towards optimal maintenance strategies. This human-guided approach ensures that the algorithm remains on the right track, continually improving its performance and ultimately leading to more efficient and cost-effective industrial maintenance practices.
Contributions. The contributions of this paper can be summarized as follows:
  • To improve the accuracy of the RUL prediction module, we adopt Transformer to address the difficulty in reflecting long-term dependencies of sequential sensor data in industrial scenarios.
@@
  • We propose a DRL-based recommendation system for determining maintenance actions, reducing both labor involvement and operational costs. Furthermore, we utilize RLHF to confirm the correct maintenance actions and improve learning efficiency.
  • A Transformer-DRL framework is proposed to optimize the prescriptive maintenance process, which can be extended to multiple manufacturing scenarios as well.
  • Extensive simulation experiment results, conducted on different datasets, demonstrate that our method outperforms baseline methods in terms of both predicting RUL and optimizing maintenance actions.
@@
Organizations. The rest of the paper is organized as follows. Section II reviews related work in the field of prescriptive maintenance utilizing machine learning algorithms. Section III formulate the problem we are trying to solve. Then, section IV describes our proposed framework for Transformer-driven prediction maintenance in detail. Next, section V presents a case study on our proposed framework using a real-world dataset. Finally, section VI concludes the paper and outlines future directions for research. Notations used across the paper is provided in Table I.
TABLE I: Notations.
Symbol Description
$\mathbf{X}$
Time series of sensor readings and operational parameters.
$R(t)$
Remaining Useful Life at time $t$: $R(t)=T-t$.
$T$
Total expected lifespan of the component.
$t$
Current operational time of the machine.
$M$
Model estimating $R(t)$ from $\mathbf{X}$.
$s$
State encapsulating health and RUL metrics.
$\mathcal{A}$
Set of maintenance actions.
$a$
Action chosen to maximize future reward.
$a^{*}$
Optimal action maximizing cumulative benefit.
$Q(s,a)$
Expected reward of action $a$ in state $s$.
$\epsilon$
Exploration rate in DRL.
$\alpha$
Learning rate in DRL.
$\gamma$
Discount factor for future rewards.
$\bm{y}$
Predicted RUL output by Transformer model.
$\bm{s}$
Observation state in DRL from RUL predictions.
$d_{\text{model}}$
Input/output dimension in Transformer.
$d_{k},d_{v}$
Dimensions of keys/queries and values in attention.
$\mathbf{Q},\mathbf{K},\mathbf{V}$
Query, key, and value matrices in attention.
$FFN(\cdot)$
Feed-forward network in Transformer.
$\bm{\Lambda}$
Attention weight matrix.
$\alpha_{ij}$
Attention weight from $j$-th input to $i$-th output.
$\bm{h}$
Hidden states in encoder.
O
Output states from decoder.
$\sigma(\cdot)$
Activation function in prediction module.
$K$
Number of machines in federated learning setup.
$f_{\text{ensemble}}$
Function for RUL prediction in multi-machine case.
$\mathbf{W},b$
Weight matrix and bias in prediction module.
PCA
Used for selecting important parts of RUL.
Figure 1: An Overview of TranDRL Framework.
@@ II Related Work
Corrective maintenance is “fault-based maintenance” and is an unplanned mode of maintenance carried out only when equipment malfunctions are detected. Preventive maintenance, or scheduled inspection and maintenance, generally involves halting the production line according to a predetermined schedule, with an emphasis on repairs before a failure occurs. These two types of maintenance methods are traditionally the most commonly adopted in the industry. However, they often lead to production losses due to their neglect of the historical operational status of the production system and components, resulting in either too frequent or too infrequent maintenance. Predictive maintenance determines the timing and methods of maintenance based on the physical model of the production system and historical monitoring data, and is therefore also referred to as condition-based maintenance. Prescriptive maintenance is generally considered an extension of predictive maintenance, optimizing maintenance strategies based on fault diagnosis results. It emphasizes identifying the inherent reasons for the occurrence of faults and subsequently directing the generation of operational and maintenance strategies [7]. Compared to traditional maintenance methods, prescriptive maintenance leverages historical system status to comprehensively forecast future system operations and optimize the resources required for maintenance, aiming to prevent waste from excessive maintenance or production losses due to malfunctions and downtime.
@@ II-A Deep Learning-Based Remaining Useful Life Prediction
In production maintenance tasks, accurately predicting the RUL of machinery has traditionally required intricate models that capture the nuances of production system operations. While many revert to statistical signal processing and machine learning to bypass comprehensive system modeling, these methods often mandate profound domain-specific knowledge, especially for tasks like signal denoising and feature extraction [8, 9]. Notably, they also entail labor-intensive manual feature engineering.
However, with machines becoming multifaceted and sensor-rich, these methods grapple with high-dimensional data. Deep learning, boosted by computational advances, emerges as a remedy. Pioneering algorithms, such as Convolutional Neural Networks (CNN) and Deep Autoencoders, have gained traction for their aptitude in fault diagnosis and predictive maintenance [10, 11]. Their intrinsic capability to distill hierarchical features from raw data eliminates manual feature engineering, fostering accurate RUL predictions.
Amongst the deep learning architectures, Transformers, originally conceived for language processing tasks [12], have recently demonstrated efficacy in time-series data processing. For instance, Zhou et al. [13] integrated a time-embedding layer, augmenting the Transformer’s time-series prediction. Yet, a full exploration of Transformers in the realm of prescriptive maintenance remains on the horizon.
@@ II-B Deep Reinforcement Learning enabled Predictive Maintenance
DRL’s concept into predictive maintenance has primarily focused on refining maintenance schedules [14, 2, 4]. While these applications exhibit DRL’s capability in mitigating machinery breakdowns, they predominantly center on forecasting malfunctions, focusing on actionable preventative measures. Gordon et al. [15] explore into employing DRL for prescriptive maintenance, but the integration of prediction and actionable recommendations remained elusive.
In existing maintenance studies, we have explored a lot at preventive and predictive methods, as mentioned in [16] and [17]. But now, there is growing interest in prescriptive maintenance, which goes beyond just predicting problems to suggesting the best actions to take. Even though the work in [16] and [17] has started to light the way, most efforts are either focused on predicting what will happen or on prescribing what to do. Even with improvements in using Transformer technology for predictions and DRL for making recommendations, there is no plan to integrate these two ideas together. Our study aims to fill this gap by introducing a way to use both the forecasting power of Transformers and the decision-making abilities of DRL to not only predict equipment failures but also suggest the best actions to prevent them.
II-C Deep Reinforcement Learning with Human Feedback
Inspired by preference theory [18], Reinforcement Learning from Human Feedback (RLHF) improves the performance of DRL by leveraging feedbacks from human, and releases the difficulty of designing reward function by incorporating human guidance and expertise [19]. To tackle the problem of informative limitation and expensive time consumption, Ibarz et al.  [20] combine expert demonstrations and trajectory preferences together and attempt to let human interact with agent directly. To enhance sample efficiency, other follow-up works leverage semi-supervised learning and data augmentation techniques into RLHF [21, 22]. Most previous works are under the non-Markov assumption which poses a threat to modeling reward function, Early et al.  [23] release this assumption and formulate the scenario into a multiple instance learning process. With the prosperity of large language model recently, RLHF also shows its power in fine-tuning step before the model development in this specific domain [24, 25]. Furthermore, in safety-critical applications like aeroengine maintenance, deviating from standard procedures can have grave consequences. Such non-standard maintenance practices might result in serious safety incidents. Combining human feedback can also help in setting safe boundaries and ensuring that the agent’s behavior remains within acceptable limits. To make the policy more reliable, the incorporation of human feedback can make the agent’s behavior more predictable and better aligned with human expectations, enhancing the explainability of its actions and building trust among users.
@@ III Problem Formulation
In the modern industrial ecosystem, machinery and equipment form the backbone of operations. However, with increased automation and reliance on these machines, any unplanned downtime can be a critical blow to productivity and financial stability. The cascade effect is not just limited to production loss; it also affects aspects like increased costs for emergency repairs, contractual penalties due to delivery delays, and even reputation damage [26]. Therefore, the primary challenge is to recommend proper maintenance actions effectively based on accurate prediction of failures of the machine.
@@ III-A Predicting Remaining Useful Life (RUL)
Predicting RUL is essential for predictive maintenance, which provides an estimation of the duration that a machine or its components will continue to function before encountering a failure. An accurate RUL prediction enable the industries to strategically schedule their maintenance activities. The RUL offers a dynamic and predictive window into the operational health of a machine. The operational data of a machinery component is typically represented as a time series $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, where each $\bm{x}_{i}$ can represent multiple sensor readings, operational parameters, or environmental conditions. The RUL at any given time $t$, denoted by $R(t)$, can be mathematically represented as: $R(t)=T-t$, where
  • $T$ is the total expected operational lifespan of the machine or component, which can vary based on its usage, environmental factors, and maintenance history.
@@
The inherent complexity lies in the fact that $T$ is not a constant but a variable that might change based on multiple factors, for example, machine conditions. The primary challenge is to develop a predictive model $M$ that can harness the rich time-series data $\mathbf{X}$ to provide an accurate approximation of $R(t)$: $M(\mathbf{X})\approx R(t)$.
III-B Optimal Maintenance Action Recommendations with DRL
While predicting the RUL offers industries a foresight into potential downtimes, it is the actionable insights derived from these predictions that truly empower industries to make informed decisions. Upon gaining prediction results of potential machine downtimes, the challenge shifts to determining the most optimal maintenance action. Leveraging the capabilities of DRL, our framework not only predicts the machinery’s future status but also suggests the best course of action to prolong its life and ensure optimal performance.
Consider a machinery component’s state $s$, which encapsulates its health, RUL, and other relevant metrics. From the state, we want to choose an action $a$ from a set of potential actions set $\mathcal{A}$ to maximize the expected future reward.
The essence of this problem is rooted in the principle of maximizing long-term rewards. We denote a value function $Q(s,a)$, which characterizes the expected cumulative benefit or reward of adopting action $a$ in state $s$. The core objective is to navigate the action space and select: $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, where
@@ III-B Optimal Maintenance Action Recommendations with DRL
  • The cumulative reward should integrate both current and weighted future’s benefits strategically.
@@ IV Proposed Framework
To solve the problem above, we propose a framework consisting of two major components: a transformer-based learning model to predict the RUL and a decision module to recommend maintenance actions as illustrated in Figure 1 and Algorithm 1. Specifically, we consider two cases for the RUL prediction, including a single machine case and a multi-machine case.
Input: Sequence of sensor readings $\mathbf{X}=(\bm{x_{1}},\bm{x_{2}},\ldots,\bm{x_{n}})$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
Output: Recommended maintenance action $a$
/* Phase 1: RUL prediction using Transformer framework */
1 Initialize the Transformer model with appropriate hyperparameters;
2 Process input sequence $\mathbf{X}$ through Transformer layers;
3 Generate RUL prediction $\bm{y}=(y_{1},y_{2},\ldots,y_{m})$;
/* Phase 2: Maintenance action recommendation using DRL */
4 Initialize DRL model with neural network weights;
5 Use RUL prediction $\bm{y}$ as the observation state $\bm{s}$;
6 Select action $a$ based on $\epsilon$-greedy policy using state $\bm{s}$;
7 Execute action $a$ and observe new state and reward;
8 Update DRL model using observed reward and new state;
9 Adjust $\epsilon$ for exploration-exploitation balance;
10 Output recommended action $a$;
return $a$
Algorithm 1 A Transformer-enabled DRL framework for maintenance selection.
IV-A Case 1: Remaining Useful Life Prediction Module for a Single Machine
Transformers, rooted in natural language processing [27], have become valuable tools in time-series data analysis. This is largely due to their proficiency in understanding complex temporal sequences. Our model leverages the Transformer’s capabilities to forecast the RUL from historical time-series sensor data. The core of this approach is the self-attention mechanism, which establishes causal relations between different time steps based on query-and-key similarities, adapting to the complexity of the data, for example, self-attention can handle complex relationships in machinery sensor data when issues depend on small changes over time in predictive maintenance.
IV-A1 Encoder Module
For a given time-series $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]$, where $n$ is the sequence length and $\bm{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$, the Encoder Module converts $\mathbf{X}$ into hidden states $\bm{h}=[h_{1},h_{2},\cdots,h_{n}]$, with each $h_{i}\in\mathbb{R}^{d_{\text{model}}}$. Each Transformer block in this module comprises:
  • Multi-Head Self-Attention Mechanism. Given matrices for query $\mathbf{Q}\in\mathbb{R}^{n\times d_{k}}$, key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$, and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$, the attention mechanism is:
    $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$
    with $d_{k}$ being the dimension of queries and keys, and $d_{v}$ the dimension of values.
  • Feed-Forward Neural Network. To capture non-linear patterns, this is represented as:
    $FFN(\bm{x}_{i})=\max(0,\bm{x}_{i}\mathbf{W}_{1}+b_{1})\mathbf{W}_{2}+b_{2},$ (1)
    with weight matrices $\mathbf{W}_{1},\mathbf{W}_{2}$ and bias terms $b_{1},b_{2}$.
IV-A2 Self-Attention Mechanism
This module facilitates the model in emphasizing or downplaying certain temporal characteristics. The attention weight $\bm{\Lambda}=[\alpha]_{i,j}$ computes the weighted sum with its $(i,j)$-th element defined as:
$\alpha_{ij}=\text{Softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{j}^{T}}{\sqrt{d_{k}}}\right),$ (2)
IV-A3 Decoder Module
After encoding the time-series data into hidden states, the decoder further refines this information to produce outputs which are conductive for RUL prediction. Given hidden states $\bm{h}^{\prime}$, the decoder module generates output states $\textbf{O}=\{\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}\}$. Each block in the decoder encompasses:
  • Multi-Head Cross-Attention Mechanism. Cross-attention takes in the encoder’s outputs and combines them with the decoder’s states. This process is represented as
    $\text{Attention}(\mathbf{Q}_{d},\mathbf{K}_{e},\mathbf{V}_{e})=\text{Softmax}\left(\frac{\mathbf{Q}_{d}\mathbf{K}_{e}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)\mathbf{V}_{e},$ (3)
    where $d_{k}$ represents the dimension of the queries and keys.
  • Feed-Forward Neural Network. This module extracts patterns in the decoder states:
    $FFN(\bm{o}_{i})=\max(0,\bm{o}_{i}\mathbf{W}_{3}+b_{3})\mathbf{W}_{4}+b_{4},$
    with weight matrices $\mathbf{W}_{3},\mathbf{W}_{4}$ and biases $b_{3},b_{4}$.
IV-A4 Prediction Module
The prediction module utilizes the states $\bm{h}$ to estimate the RUL. This involves one or several dense layers, which can be expressed as
$\text{RUL}=\sigma(\bm{h}\mathbf{W}+b),$
where the weight matrix $\mathbf{W}$, bias $b$, and activation function $\sigma$ ensure that the RUL remains within the range.
Input: Input sequence $\mathbf{X}=(\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{n})$